OpenStudy (anonymous):

A cornmeal feed is 10% protein, and a feed supplement is 40% protein. How much of each must be combined to get 2400 pounds of feed that is 19% protein?

OpenStudy (anonymous):

Let c and p be the final weights of the cornmeal feed and the feed supplement to be combined, respectively. The sum of these weights of have to equal 2400. The sum of the protein weight contributed by c and p has to be 19% of 2400 lbs. Solve the following for c and p:\[\left\{c+p=2400,\frac{10}{100}c+\frac{40}{100}p=\frac{19}{100}2400\right\}\]c = 1680 lbs and p = 720 lbs.